OpenStudy (anonymous):

could you possibly help with this one. find the exact value of the expression cos pi/3+ sin pi/2

OpenStudy (anonymous):

\[\cos(\alpha +\beta)=\cos(\alpha)\cos(\beta)-\sin(\alpha )\sin(\beta)\] use \[\cos(\frac{\pi}{3})=\frac{1}{2}, \cos(\frac{\pi}{2})=0,\sin(\frac{\pi}{3})=\frac{\sqrt{3}}{2}, \sin(\frac{\pi}{2})=1\]